The equivalent resistance of two resistors R1 and R2 is given by π‘…π‘’π‘ž = 𝑅1 + 𝑅2 and π‘…π‘’π‘ž =



𝑅1𝑅2



𝑅1+𝑅2



when they are connected in series and parallel respectively. If 𝑅1 = 22.0 Ξ© Β±



5% π‘Žπ‘›π‘‘ 𝑅2 = 47.0 Β± 5% , find the maximum error in the calculation of their equivalent



resistance when they are connected in series and parallel.

1
Expert's answer
2021-11-15T17:50:35-0500

"R_s=R_1+R_2"

Now errors

βˆ†Rs =(βˆ†R1+βˆ†R2 )

"∆R_s=\pm({0.05}+{0.05})=\pm0.10"

"R'_s=R_1+R_2=22+48=70\Omega"

Now

"R_{eq}=70\pm0.10"

"\frac{1}{R_{eq}}=\frac{1}{R_1}+\frac{1}{R_2}"

"\frac{∆R_{eq}}{R_{eq}}=\frac{∆R_1}{R_1}+\frac{∆R_2}{R_2}"

"\frac{∆R_{eq}}{R_{eq}}=\pm(\frac{0.05}{22}+\frac{0.05}{47})"


"\frac{∆R_{eq}}{R_{eq}}=\pm(0.00227+0.00106)=\pm(0.00303)"

"R_{eq}=\frac{R_1R_2}{R_1+R_2}"

"R_{eq}=\frac{22\times47}{22+47}=14.98\Omega"


"∆R_{eq}=14.98\times0.00303=0.04988\Omega"

"R_{eq}=14.98\pm0.04988\Omega"